find the equation of the line that is parallel to 5x+2y=-1 and contains the point (-4,-2)
Answer by lwsshak3(11628)

ind the equation of the line that is parallel to 5x+2y=-1 and contains the point (-4,-2)
**
5x+2y=-1
rewrite in standard form for a straight line
2y=-5x-1
y=-5x/2-1/2
slope=-5/2
..
Slope of line parallel to given line=-5/2
Equation: y=-5x/2+b
solving for b using given point (-4,-2)
-2=(-5*-4)/2+b
-2=10+b
b=-12
Equation of parallel line:
y=-5x/2-12